Tags:
create new tag
view all tags

Shift Log 20120418

SB LN

09:00

09:30

  • Checked lens in tunnel. No damage from previous shift.

10:00 Machine recovered, steer and quad settings

steer_20120418.png quad_20120418.png

OTR

  • Data directory 20120418-shift
  • School boy.... wink Move OTR in ...SB didn't notice...
  • Saved positions as moving OTR 20120418_1015_pos.dat
  • Set OTR filter : 3

  • Scans set to minimum Alar OTR size
    • QM15FF scan (Okugi set : -38)
    • QM14FF scan (Okugi set : -5.425, set to -13.425)
    • QM15FF rescan (Set to -38.000)
    • QS scan (Set to nominal -0.495)
    • QK4 scan (Set to 2.610)
    • Polariser scan

Screen pos : 8474 Laser location as before

Alignment

_1331 Vertical scan, 0 to 500, set to 250
Removed horizontal orbit offset within 100 um, 100 urad of zero
_1356 Horizontal scan, 0 to 500, set to 275
_1404 Vertical scan, 0 to 500, inconclusive. _14XX Vertical scan, 0 to 700, set to 600.

Horizontal alignment Vertical alignment
20120418_1410_lws_xy_errbar_plot.png 20120418_1356_lws_xy_errbar_plot.png

  • Removed screen, LW detector on. Background high. Detector saturated with 10dB attenuation - 8k previously.

XX:XX Okugi LW only optics

Mag Current Old shift
QM16FF -37.211 8.378
QM15FF 41.927 21.037
QM14FF -101.892 -88.730
QM13FF 37.114 39.040
QM12FF -28.423 -21.000
QM11FF 9.560 1.000

ZH1FF ZH10X Comment
-0.289A -0.004 Oribt good, background 12k @ 10db

Can use these two to reduce background to around 9000 but have 1mm horizontal offset at LWIP as well as 600urad angle. Trying magnet movers of upstream magnets to attempt to compensate. Unfortunately, QM11FF QD10BFF movers are broken and so as they're horribly misplaced, they give the beam a huge kick. Have to offset with others afterwards, but probably generating a lot of background.

Changed Values of QM13 and QM12 (after us in that order).

QM13FF QM12FF Comment Background
36.622A -11.802A Before 13k
27.622A -9.802A After 7k

Terunum-san is maintaining excellent charge with 1.2x10^10. Reduce laser intensity to compare background levels.

_1440_log with laser intensity at 35% - around 1.2x10^10 in extraction line - background around 5k _1443_log with laser intensity at 15% - around 0.6x10^10 in extraction line - background around 2.7k

Returned to higher charge. Background level better than before.

  • Timing using old scope. New scope - dodgy representation of timing. Set phase to 4.0V
  • Found collisions at 4pm at approximately CV:674um

  • Phase Scan to optimise. _1552_lws.dat Quite jumpy - phasing problem

  • Initial Phase Scan:
20120418_1558_lws_xy_fit_plot.png

  • Changed main LW TD2 from 1538 (base 10 on EDM) to +1 - phase scan ) _1603 - right on edge at 7.5V
  • Return to 4V, TD2 set to 1538 and long range scan 0 - 10V _1608_lws.dat
  • TD2 set to 1537 and long range scan from 0 - 10V _1613_lws.dat
  • TD2 set to 1536 and long range scan from 0 - 10V _1617_lws.dat

  • 100 pulse log during 4:30pm meeting to investigate phase stability on peak. _1624_log.dat

  • Horizontal scan over 1mm from -600 to 600 failed to encompass beam horizontal - pretty big indeed - OTR measurement 200um

  • Horizontal Measurement:
20120418_1702_lws_xy_errbar_plot.png

Noticed timing instability where signal would modulate signal by several thousand ADC counts. Investigated and thought it could be 357MHz phasing with respect to beam. After about an hour of tracing signals and investigating, found it to be the oscillator. Although it was locked, it had sidebands on the spectrum analyser. Didn't tune cavity, but merely cleaned it. Output power was 275mW before and 350mW after.

  • Laser locked cleanly, collisions found at 1538 (+1) vs before. Phase max is now around 9V. Tried to find collisions in next RF bucket, but without success. Ok at 9V.
  • Phase scan to optimise. Set at 9.5V _1901_lws.dat
  • Vertical scan _1907_lws.dat

Changing QM15FF

QM15FF (A) Filename $\sigma_{y}$ ($\mu$m)
-38.0 _1907_lws.dat 7.5
-36.0 _1918_lws.dat 8.0
-37.0 _1926_lws.dat 8.0
-38.0 _1933_lws.dat 58.0
-38.0 _1941_lws.dat 14.0
-39.0 _1951_lws.dat 36.0
-40.0 _1958_lws.dat 12.6
-40.0 _2006_lws.dat 11.5

Fit results are fairly meaningless. Most fits have a secondary peak to the right on most plots. The beam is more than likely rolled and in addition plots of the ipoffset in the horizontal show it wandering by up to 100$\mu$m. If the electron beam is moving as the scan progresses vertically, we would expect a secondary scan, which is what we see.

  • Unusual Vertical Profile:
20120418_1933_lws_xy_errbar_plot.png

  • Aspect ratio of the beam is quite large.

  • Perform skew quad scan and then change optics.
  • QM15FF returned to -38.0A

QK4X (A) Filename $\sigma_{y}$ ($\mu$m)
2.61 _2014_lws.dat 9.2
-2.39 _2022_lws.dat 13.5
-7.39 _2029_lws.dat 48.2
-12.39 _2038_lws.dat didn't fit
-17.39 _2045_lws.dat 6.2
7.61 _2055_lws.dat 10.6
12.61 _2103_lws.dat 12.0
17.61 _2110_lws.dat 7.8
-2.39 _2121_lws.dat 11.74
-7.39 _2130_lws.dat  

20120418_2029_lws_xy_fit_plot.png 20120418_2045_lws_xy_fit_plot1.png

Switching to Okugi-san's unmatched optics, which should give us a smaller size. Saved set file with previous Okugi-san matched optics - set12apr182147.dat.

XX:XX Okugi LW only optics

Mag Current Old shift
QM16FF -37.211 8.378
QM15FF 41.927 21.037
QM14FF -101.892 -88.730
QM13FF 37.114 39.040
QM12FF -28.423 -21.000
QM11FF 9.560 1.000

11pm

Loaded Okugi-san's new unmatched optics with smaller beam size. AA reduced background from saturation to around 4k. Need to include magnet movers for absolute positioning. Moved magnet movers so although Bpms say we're quite similar, screen alignment says we're actually around 1mm out.

  • New alignment CV: 1200um, CH: 450um.

Moved to old magnet mover location

  • QM14FF : -50, 50
  • QM13FF : -300, 300

Tuned background again using ZH1FF, ZV1FF, ZH1X

Loaded new Okugi-san optics. Minimum beam size vertically 7.22 +- 0.26 um.

  • Reloaded optics from shift last week ('old shift') in table above. Beam initially seems smaller on OTR.

  • Tweaked QM14FF to -86.73A Looks very clear minima in two lobe distribution.
  • Set file 'set12apr19_0058.dat
  • OTR vertical measurement 2.04 +- 0.51 um.

Edit | Attach | Watch | Print version | History: r10 < r9 < r8 < r7 < r6 | Backlinks | Raw View | Raw edit | More topic actions

Physics WebpagesRHUL WebpagesCampus Connect • Royal Holloway, University of London, Egham, Surrey TW20 0EX; Tel/Fax +44 (0)1784 434455/437520

Topic revision: r10 - 24 Oct 2013 - LaurieNevay

 
This site is powered by the TWiki collaboration platform Powered by PerlCopyright © 2008-2023 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding RHUL Physics Department TWiki? Send feedback